OpenStudy (anonymous):

Given the function f(x) = 2(x + 6), find x if f(x) = 22. 56 5 38 17

jimthompson5910 (jim_thompson5910):

f(x) = 2(x + 6) 22 = 2(x + 6) 22 = 2x + 2*6 22 = 2x + 12 22-12 = 2x + 12-12 10 = 2x 10/2 = 2x/2 5 = x x = 5

OpenStudy (anonymous):

x=5